A sound wave has a frequency of 680 Hz in air and a wavelength of 0.48 m. What is the temperature of the air? Assume the velocity of sound at 0◦C is 330 m/s. Answer in units of ◦C.
Answer #1

Suppose temperature of the air = T

f = frequency of sound wave at this temperature T = 680 Hz

wavelength = 0.48 m

v = velocity of sound wave at this temperature T = f * wavelength = 680 * 0.48 = 326.4 m/s

vo = velocity of sound at 0◦C = 330 m/s

formula:   v = vo + 0.61 *T

326.4 = 330 + 0.61 *T

T = - 5.9 degree celcius